By considering the summation: 2+22+23+…..+2n-1

A). What is the formula used for the summation? ________________________

B). Prove that your formula is correct for all integers n>=2 by mathematical induction.
Answer #1

A) For every natural number n, 21 + 22 + 23... + 2n-1 = 2n − 2 (Lets call this equation P(n))

B) Proof:

  • Base case:

Let n = 2, then

LHS

=====

= 2

RHS

=========

= 22 - 2 = 4 - 2= 2

  • Inductive Hypothesis:

Let us assume P(k) is true for n = k.

  • Inductive Step:

We need to prove that the equation P(k+1) is true for n = k+1

For P(k+1):

LHS

===========

= 21 + 22 + 23 + .. + 2k-1 + 2k = (21 + 22 + 23 + .. + 2k-1) + 2k

=  2k− 2 + 2n

= 2.2k - 2 = 2k+1 - 2 (proved)